### v1

@@ 3.2 New Reduction: Kaczmarz Method and Stepwise-Optimal Stochastic Gradient Descent
It is known that when each task contains just one row, each update in the Kaczmarz method corresponds to a gradient step on with a specific “normalizing” step size (Needell et al., 2014). That is, noticing that in rank-$1$ we have $\mathcal{L}_{\tau(t)}(\mathbf{w})=\frac{1}{2}\big{\|}{\mathbf{x}_{\tau({t})}^{\top}\mathbf{w}-y_{\tau({t})}}\big{\|}^{2}$, the Kaczmarz updates hold
$\mathbf{w}_{t}=\mathbf{w}_{t-1}-\frac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\big{(}{\mathbf{x}_{\tau(t)}^{\top}\mathbf{w}_{t-1}-y_{\tau(t)}}\big{)}\mathbf{x}_{\tau(t)}=\mathbf{w}_{t-1}-\frac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\nabla_{\mathbf{w}}\mathcal{L}_{\tau(t)}(\mathbf{w}_{t-1})\,.$ (1)
@@
In the realizable case (Assumption 1) under any ordering $\tau$, the block Kaczmarz method is equivalent to SGD with a step size of $\eta=1$, applied w.r.t. a convex, $1$-smooth least squares objective: $\big{\{}f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}\big{\}}_{m=1}^{T}$. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 2 and 3 coincide.
The reduction follows from the next lemma, which states the convexity and smoothness of $f_{m}$ and expresses the gradient $\nabla_{\mathbf{w}}f_{\tau(t)}$, subsequently substituted into $\big{(}{\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})}\big{)}$ to complete the proof.
Lemma 6 (Properties of the Modified Objective).
Consider any realizable task collection such that ${\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m}},\forall m\in\left[{T}\right]$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
@@ 4 Rates for Random-Order Continual Linear Regression and Kaczmarz
This section focuses on improving the best upper bound known in prior continual learning literature on random orderings, summarized in Table 1. Specifically, for with-replacement random orderings, Evron et al. (2022) proved a forgetting rate of $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]=\mathcal{O}\!\left({\frac{d-\bar{r}}{k}}\right)$ where $\bar{r}\triangleq\frac{1}{T}\sum_{m}\operatorname{rank}(\mathbf{X}_{m})$. Notably, this rate depends on the problem dimensionality $d$, raising concerns when generalizing insights from linear models to deep neural networks—which are often highly overparameterized (e.g., in the NTK regime). Encouragingly, that paper only provided (implicitly) a $1/k$ lower bound for the worst-case forgetting, calling for further research into narrowing this gap.
@@
Table 1: Forgetting and Loss Rates in Continual Linear Regression (and Block Kaczmarz). Upper bounds apply to any $T$ realizable tasks (or blocks). Lower bounds indicate worst cases, i.e., specific constructions. Random ordering bounds apply to the expected forgetting (or loss).
We omit mild constant multiplicative factors and an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ term. Finally, ${a\!\wedge\!b\triangleq\min(a,b)}$. Recall: $k=\,$iterations; $d=\,$dimensionality; $\bar{r},r_{\max}=\,$average and maximum data matrix ranks.
@@
Evron et al. (2022) Upper $\frac{d-\bar{r}}{k}$ $\frac{T^{2}}{\sqrt{k}}\!\wedge\!\frac{T^{2}(d\!-\!r_{\max})}{k}$
Kong et al. (2023) Upper $\frac{T^{3}}{k}$
@@
Evron et al. (2022) Lower $\frac{1}{k}$ (*) $\frac{1}{T}$ (*) $\frac{T^{2}}{k}$
@@ 4.1 A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
Theorem 7 (Parameter-Dependent Forgetting Rate for Random With Replacement).
@@
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq\frac{\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2e(k-1)}\,,\quad\quad\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]\leq\frac{3\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2\left({k-2}\right)}\,,$
@@
The $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ factor is a scaling term, generally unavoidable in this area. The proof is outlined below, and detailed in Appendix C. Our analysis is related to a recent work (Guo et al., 2022) that also analyzes a linear map to characterize the weak error (resembling our loss). They achieve polynomial rates, that, unlike ours, involve matrix properties related to the condition number.
@@ 4.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
Theorem 8 (Universal Forgetting Rate for With-Replacement Random Ordering).
@@
We prove this result in Section D.1, by leveraging the connections between continual learning and SGD. Specifically, in Section 3 we showed that continual linear regression and the Kaczmarz method, are equivalent to SGD with a step size of exactly $1$ on a related least squares problem. Our result follows from our novel last-iterate SGD bounds that, crucially, apply even to that specific step size. To ease readability, here we focused on a continual learning perspective, deferring last-iterate analysis to Section 5.
@@ 4.3 Random Task Orderings Without Replacement
Evron et al. (2022) suggested defining forgetting as “catastrophic” only when $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$.They presented such an adversarial case with a deterministic task ordering where $k=T\to\infty$. They showed that task recurrence, under cyclic or random orderings, mitigates forgetting. So far, in random orderings, it was hard to isolate the effect of randomness from that of repetitions. It was thus unclear whether catastrophic forgetting can be alleviated by randomly permuting the tasks. Below, we provide the first result demonstrating that no recurrence is needed under random orderings.
Theorem 9 (Forgetting Rates for Without-Replacement Random Ordering).
@@
$\mathop{\mathbb{E}}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right],~{}\mathop{\mathbb{E}}_{\tau}\left[F_{\tau}(k)\right]\leq\min\left(\frac{7}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}+1}{k-1}\right){\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}.$
The proof of the dimensionality-dependent term is similar to the one of the with-replacement case, given in Section D.1.2 of Evron et al. (2022), but requires a more careful upper bound on the (in-sample) forgetting. The proof of the dimensionality-independent term again relies on last-iterate analysis, as presented in Section 5.2. Both proofs are given in Section D.2.
In Appendix A, we discuss relations between without-replacement random orderings and other areas like shuffle SGD.
@@ 5 Last-Iterate SGD Bounds for Linear Regression
In this self-contained section, we derive last-iterate guarantees for SGD in the realizable stochastic least squares setup. Motivated by the connection with continual regression discussed in Section 3, we focus on regression problems that are $\beta$-smooth individually, and obtain upper bounds for the last SGD iterate that apply for a significantly wider range of step sizes compared to prior art (Varre et al., 2021). Notably, this is the first time convergence of SGD in this setup is established for a range of step sizes completely independent of the optimization horizon. Table 2 compares our bounds with related work and classic results.
Many recent works study SGD in the realizable least squares setting (Ge et al., 2019; Vaswani et al., 2019; Berthier et al., 2020; Zou et al., 2021; Varre et al., 2021; Wu et al., 2022) under a slightly more general noise model. These studies are primarily motivated by the connection between this setup and deep neural networks in the overparameterized regime (Ma et al., 2018), where the model is expressive enough to perfectly fit the training data. With the exception of Varre et al. (2021), most of these works focus on non-fixed step size schedules and/or provide guarantees for the average iterate. Further discussion of related work can be found in Appendix A. Proceeding, throughout this section we consider stochastic, realizable least squares problem as defined next.
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n_{i}\times d},\mathbf{b}_{i}\in\mathbb{R}^{n_{i}},~{}\forall i\in\mathcal{I}$. We specifically focus on $\beta$-smooth functions, i.e., $\left\|{\mathbf{A}_{i}}\right\|\leq\beta,\forall i\in\mathcal{I}$, under a realizable assumption, i.e., $\exists\mathbf{w}_{\star}\in\mathbb{R}^{d}:\bar{f}(\mathbf{w}_{\star})=0$.
Table 2: State-of-the-art Loss Bounds for Fixed-Step-Size SGD. We consider stochastic convex optimization with an objective $\bar{f}(\mathbf{w})\!\triangleq\!\mathbb{E}_{\xi}f(\mathbf{w};\xi)$, where $f(\cdot;\xi)$ is $\beta$-smooth almost surely, $\sigma^{2}\geq\mathbb{E}\|\nabla f(\mathbf{w};\xi)-\nabla\bar{f}(\mathbf{w})\|^{2}$, $\sigma_{\star}^{2}\triangleq\mathbb{E}\|\nabla f(\mathbf{w}_{\star};\xi)-\nabla\bar{f}(\mathbf{w}_{\star})\|^{2}$, and $G>0$ is such that $\|\nabla f(\mathbf{w};\xi)\|\leq G$ for any $\mathbf{w}$ and $\xi$. Dependence on constant numerical factors and the distance to an optimal solution is suppressed.
Stochastic (*) Shamir and Zhang (2013) $\frac{1}{\eta T}+\eta G^{2}\log T$
Deterministic Smooth ($\sigma=0$) Nesterov (1998) $\frac{1}{(2-\eta\beta)\eta T}$
Stochastic Smooth Lan (2012) $\frac{1}{\eta T}+\eta\sigma^{2}$
Liu and Zhou (2024) $\frac{1}{\eta T}+\eta\sigma^{2}\log T$
Stochastic Smooth Realizable ($\sigma_{\star}=0$) Srebro et al. (2010) $\frac{1}{(1-\eta\beta)\eta T}$
Stochastic Regression Realizable ($\sigma_{\star}=0$) Bach and Moulines (2013) $\frac{1}{\eta T}$
Varre et al. (2021) $\frac{1}{(1-2\eta\beta\log T)\eta T}$
Ours (2025) $\frac{1}{(2-\eta\beta)\eta T^{1-\eta\beta(1-\eta\beta/4)}}$
(*) Shamir and Zhang (2013) consider bounded domains; Orabona (2020); Liu and Zhou (2024) obtain similar bounds for the unconstrained case. For non-fixed step sizes Jain et al. (2019) obtain minimax optimal bounds without log factors.
@@
Theorem 10 (Last-Iterate Bound for Realizable Regression).
@@
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular, for $\eta=\frac{1}{\beta}$:
$\mathbb{E}\bar{f}(\mathbf{w}_{T})\leq\frac{e\beta D^{2}}{2\sqrt[4]{T}}.$
The important feature of Theorem 10 is the $(2-\eta\beta)$ factor in the denominator, replacing the common $(1-\eta\beta)$ of the standard analysis. This difference makes our theorem applicable to the continual regression setting which requires setting $\eta=1/\beta$ (Reduction 2). In addition, for $\eta={1}/{\left({\beta\log T}\right)}$, we recover the near-optimal rate obtained by Varre et al. (2021), i.e., $\mathbb{E}\bar{f}(\mathbf{w}_{T})=\mathcal{O}\left({\frac{\beta D^{2}\log T}{T}}\right).$
5.1 Analysis overview
In this section, we briefly outline the proof of Theorem 10, which follows immediately by combining the two lemmas below (while noting that $\eta<2/\beta\Rightarrow e^{\eta\beta\left(1-\eta\beta/4\right)}\leq e$). The first step of the proof is to establish a regret bound for SGD when applied to $f(\mathbf{w};i_{1})\ldots f(\mathbf{w};i_{T})$, holding for any step size $\eta<2/\beta$. This already departs from the standard $\eta<1/\beta$ mandated by standard analysis. All proofs for this section are given in Section E.1.
Lemma 11 (Gradient Descent Regret Bound for Smooth Optimization).
@@ 5.1 Analysis overview
With the above lemma in place, the second and main step of the analysis is to relate the loss of the last SGD iterate to the regret of the algorithm. For this, we carefully adapt an existing approach for last-iterate convergence in the non-smooth case (Shamir and Zhang, 2013). The result, given below, is slightly more general to accommodate without-replacement sampling, which we address in the next section.
Lemma 12.
@@
5.2 Extension to Without-Replacement SGD
Here, we extend Theorem 10 to the without-replacement setting. Specifically, we assume =$\mathcal{I}=\{0,\ldots,n-1\}$, $\mathcal{D}=\textrm{Unif}(\mathcal{I})$, and consider gradient descent w.r.t. a uniformly random permutation of the training examples. For any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, step size $\eta>0$, and $\pi_{t}\sim\textrm{Unif}(\mathcal{I})$ sampled without replacement:
Theorem 13 (Without-Replacement Bound).
Consider the $\beta$-smooth, realizable Setup 1. Define for all $T\!\geq\!2$, $\hat{f}_{0:T}(\mathbf{w})\!\triangleq\!\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};\pi_{t})$. Then, without-replacement SGD (Eq. (3)) with step-size $\eta<{2}/{\beta}$, holds:
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular,
$\eta=\frac{1}{\beta\log T}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq{\displaystyle\frac{14\beta D^{2}\log T}{T}}\,,\quad\quad\eta=\frac{1}{\beta}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq\frac{7\beta D^{2}}{\sqrt[4]{T}}\,.$
The proof, given in Section E.2, is based on the algorithmic stability of SGD (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010; Hardt et al., 2016), and more specifically, on a variant of stability, suitable for without replacement sampling (Sherman et al., 2021; Koren et al., 2022).

### v2

@@ 3.2 New Reduction: Kaczmarz Method and Stepwise-Optimal Stochastic Gradient Descent
It is known that when each task contains just one row, each update in the Kaczmarz method corresponds to a gradient step on with a specific “normalizing” step size [51]. That is, since in rank-$1$ we have $\mathcal{L}_{\tau(t)}(\mathbf{w})=\frac{1}{2}\big{\|}{\mathbf{x}_{\tau({t})}^{\top}\mathbf{w}-y_{\tau({t})}}\big{\|}^{2}$, the Kaczmarz updates hold
$\mathbf{w}_{t}=\mathbf{w}_{t-1}-\tfrac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\big{(}{\mathbf{x}_{\tau(t)}^{\top}\mathbf{w}_{t-1}-y_{\tau(t)}}\big{)}\mathbf{x}_{\tau(t)}=\mathbf{w}_{t-1}-\tfrac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\nabla_{\mathbf{w}}\mathcal{L}_{\tau(t)}(\mathbf{w}_{t-1})\,.$ (1)
@@
In the realizable case (Assumption 2.2) under any ordering $\tau$, the block Kaczmarz method is equivalent to SGD with a step size of $\eta=1$, applied w.r.t. a convex, $1$-smooth least squares objective: $\big{\{}f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right)}\right\|^{2}\big{\}}_{m=1}^{T}$. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 2 and 3 coincide.
The reduction follows from the next lemma, which states the convexity and smoothness of $f_{m}$ and expresses the gradient $\nabla_{\mathbf{w}}f_{\tau(t)}$, subsequently substituted into $\big{(}{\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})}\big{)}$ to complete the proof. The proof for the lemma is given in App. B.
Lemma 3.1 (Properties of the Modified Objective).
Consider any realizable task collection such that ${\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m}},\forall m\in\left[{T}\right]$. Define ${f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\left({\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right)}\right\|^{2}}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
@@
    Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}(\mathbf{w})$
@@
    Gradient:       $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$
@@ 4 Rates for Random-Order Continual Linear Regression and Kaczmarz
This section focuses on improving the best upper bound known in prior continual learning literature on random orderings, summarized in Table 1. Specifically, for with-replacement random orderings, Evron et al. [15] proved a forgetting rate of $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]=\mathcal{O}\!\left({\frac{d-\bar{r}}{k}}\right)$ where $\bar{r}\triangleq\frac{1}{T}\sum_{m}\operatorname{rank}(\mathbf{X}_{m})$. Notably, this rate depends on the problem dimensionality $d$, raising concerns when generalizing insights from linear models to deep networks, which are often heavily overparameterized (e.g., in the NTK regime). Encouragingly, that paper only provided (implicitly) a $1/k$ lower bound for the worst-case forgetting, calling for further research into narrowing this gap.
@@
Table 1: Forgetting and Loss Rates in Continual Linear Regression (and Block Kaczmarz). Upper bounds apply to any $T$ realizable tasks (or blocks). Lower bounds indicate worst cases, i.e., specific constructions. Random ordering bounds apply to the expected forgetting (or loss). We omit mild constant multiplicative factors and an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2\!}R^{2}$ term. Finally, ${a\wedge b\triangleq\min(a,b)}$. Recall: $k=\,$iterations; $d=\,$dimensionality; $\bar{r},r_{\max}=\,$average and maximum data matrix ranks.
@@
Evron et al. [15] Upper $\frac{d-\bar{r}}{k}$  $\frac{T^{2}}{\sqrt{k}}\wedge\frac{T^{2}(d-r_{\max})}{k}$
Kong et al. [34] Upper $\frac{T^{3}}{k}$
@@
Evron et al. [15] Lower $\frac{1}{k}$ (*) $\frac{1}{T}$ (*) $\frac{T^{2}}{k}$
@@ 4.1 A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
Theorem 4.1 (Parameter-Dependent Forgetting Rate for Random With Replacement).
@@
$\mathbb{E}_{\tau}\!\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq\frac{\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2e(k-1)},\,\,\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]\leq\frac{3\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2\left({k-2}\right)},$
@@
The $\left\|{\mathbf{w}_{\star}}\right\|^{2}\!R^{2}$ scaling term is generally unavoidable. Our proof, given in App. C, is related to a recent work [23] that characterized the weak error (resembling our loss) by analyzing a linear map. Unlike ours, the polynomial rates they derive involve matrix properties related to the condition number.
@@ 4.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
Theorem 4.2 (Universal Forgetting Rate for With-Replacement Random Ordering).
@@
We prove this result in Section D.1 by leveraging the connections between CL and SGD. Specifically, in Section 3 we showed that continual linear regression is equivalent to SGD with a step size of exactly $1$on a related least squares objective that bounds the original continual learning loss. Our result then follows from our novel last-iterate SGD bounds that, crucially, apply even to that specific step size. To ease readability, here we focused on a CL perspective, deferring last-iterate analysis to Section 5.
@@ 4.3 Random Task Orderings Without Replacement
Evron et al. [15] suggested defining forgetting as “catastrophic” only when $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$.They presented such an adversarial case with a deterministic task ordering where $k=T\to\infty$, and showed that task recurrence, under cyclic or random orderings, mitigates forgetting. So far, in random orderings, it was hard to isolate the effect of randomness from that of repetitions. It was thus unclear whether catastrophic forgetting can be alleviated by randomly permuting the tasks. Below, we provide the first result demonstrating that no recurrence is needed under random orderings.
Theorem 4.3 (Forgetting Rates for Without-Replacement Random Ordering).
@@
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right],~{}\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]\leq\min\left(\frac{7}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}+1}{k-1}\right){\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}.$
The proof of the dimensionality-dependent term is similar to the one of the with-replacement case, given in Section D.1.2 of Evron et al. [15], but requires a more careful upper bound on the (in-sample) forgetting. The proof of the dimensionality-independent term again relies on last-iterate analysis, as presented in Section E.2. Both proofs are given in Section D.2.
App. A discusses connections between our result above and related areas like shuffle SGD.
@@ 5 Last-Iterate SGD Bounds for Linear Regression
In this self-contained section, we derive last-iterate guarantees for SGD in the realizable stochastic least squares setup. Motivated by the connection with continual regression discussed in Section 3, we focus on regression problems that are $\beta$-smooth individually, and obtain upper bounds for the last SGD iterate that apply for a significantly wider range of step sizes compared to prior art [67]. Notably, this is the first time convergence of SGD in this setup is established for a range of step sizes completely independent of the optimization horizon. Table 3 in App. A compares our bounds with related work and classical results in the field.
Recent work has analyzed SGD in the realizable least squares setting, with most assuming a somewhat more general noise model [18, 68, 4, 73, 67, 69]. These studies are primarily motivated by connections to deep networks in the overparameterized regime [45], where models are expressive enough to perfectly fit the training data. With the exception of Varre et al. [67], most of these works focus on non-fixed step sizes and/or provide guarantees for the average iterate. See App. A for further discussion. Here, we study a stochastic, jointly realizable least squares problem, as defined next.
@@
where $\mathbf{A}_{i}\in\mathbb{R}^{n_{i}\times d},\mathbf{b}_{i}\in\mathbb{R}^{n_{i}},~{}\forall i\in\mathcal{I}$. We specifically focus on $\beta$-smooth functions, i.e., ${\left\|{\mathbf{A}_{i}^{\top}\mathbf{A}_{i}}\right\|\leq\beta},\forall{i\in\mathcal{I}}$, under a realizable assumption, i.e., $\exists\mathbf{w}_{\star}\in\mathbb{R}^{d}:\bar{f}(\mathbf{w}_{\star})=0$.
@@
Theorem 5.1 (Last-Iterate Bound for Realizable Regression With Replacement).
@@
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular, for $\eta=\frac{1}{\beta}$, $\mathbb{E}\bar{f}(\mathbf{w}_{T})\leq\frac{e\beta D^{2}}{2\sqrt[4]{T}}$.
The important feature of Theorem 5.1 is the $(2-\eta\beta)$ factor in the denominator, replacing the common $(1-\eta\beta)$ of the standard analysis. This difference makes our theorem applicable to the continual regression setting which requires setting $\eta=1/\beta$ (Reduction 2). In addition, for $\eta={1}/{\left({\beta\log T}\right)}$, we recover the near-optimal rate obtained by Varre et al. [67], i.e., $\mathbb{E}\bar{f}(\mathbf{w}_{T})=\mathcal{O}\left({\frac{\beta D^{2}\log T}{T}}\right).$
Extension to Without-Replacement SGD.
In Section E.2, we extend Theorem 5.1 to the setting of SGD without replacement. The proof leverages algorithmic stability for SGD [6, 61, 25], focusing on a variant tailored to without-replacement sampling [63, 35]. In particular, we establish a new bound for this variant in the smooth and realizable regime, which has not appeared in prior work.
Analysis Overview.
Here, we briefly outline the proof of Theorem 5.1, which follows immediately by combining the two lemmas below (while noting that $\eta<2/\beta\Rightarrow e^{\eta\beta\left(1-\eta\beta/4\right)}\leq e$). The first step of the proof is to establish a regret bound for SGD when applied to $f(\mathbf{w};i_{1})\ldots f(\mathbf{w};i_{T})$, holding for any step size $\eta<2/\beta$. This already departs from the standard $\eta<1/\beta$ mandated by standard analysis. All proofs for this section are given in Section E.1.
Lemma 5.2 (Gradient Descent Regret Bound for Smooth Optimization).
@@
The second and main step of the analysis is to relate the loss of the last SGD iterate to the regret of the algorithm. For this, we carefully adapt an existing approach for last-iterate convergence in the non-smooth case [62]. The result, given below, is slightly more general to accommodate without-replacement sampling, addressed in the next section.
Lemma 5.3.
@@
$\mathbb{E}f(\mathbf{w}_{T},i_{T})\leq(eT)^{\eta\beta\left(1-\eta\beta/4\right)}\mathbb{E}\left[\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w}_{t};i_{t})\right],$